The tens digit is 5 more than the unit digit. If the sum of their squares is 53, find the number.
Answer by ewatrrr(24785)

 
Hi,
The tens digit is 5 more than the unit digit
let x and (x+5) represent the unit's and ten's digits respectively
Question states***
x^2 + (x+5)^2 = 53
2x^2 + 10x - 28 = 0
x^2 + 5x - 14 = 0
factoring
(x+7)(x-2) = 0
(x+7)= 0 x = -7 tossing out negative number for a digit amount
(x-2) = 0 x = 2, one's digit, ten digit is 7. Number is 72.
CHECKING our Answer***
2^2 + 7^2 = 4 + 49 = 53